OpenStudy (anonymous):

Please help simplify. (5 ab^2) 2 ------- ( C^3 )

OpenStudy (diyadiya):

\[\frac{25a^2b^4}{c^3}\]

OpenStudy (anonymous):

that must've looked more intimidating than it was--i GOT that! thank you so much tho!

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

care to try a slightly different one?...i could you a walkthru if you can...

OpenStudy (diyadiya):

be right back!

OpenStudy (anonymous):

6) Write with positive exponents: 6^-2 a^-3^b4 ------------- c^-4 d^-2

OpenStudy (anonymous):

how do i convert this?

OpenStudy (diyadiya):

\[a^{-m}= \frac{1}{a^m}\]

OpenStudy (anonymous):

i don't understand...

OpenStudy (diyadiya):

so \[a^{-3} = \frac{1}{a^3}\] do the same for the rest

OpenStudy (anonymous):

wait--i DO get it!

OpenStudy (diyadiya):

??

OpenStudy (diyadiya):

\[\huge\frac{6^{-2}a^{-3}b^4}{c^{-4}d^{-2}}\] is this your question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the question is to convert 3 into positive exponents